6. (10 pts.) Assume you are the manager of Faygo (Michigan’s own soft drink!), where the monthly demand for Faygo at the local gas station is QD = 600 – 800P and the supply function (including sales tax) is QS = 600P – 100.

  1. Solve for equilibrium price and quantity at the local gas station.
  1. After seeing the most recent health statistics, Michigan’s congress realizes Michigan residents have a problem with obesity. Wanting to do their part to fight obesity, congress decides their best option is to place a tax on sugary drinks. They decide to impose an excise tax of $0.10 on each soft drink sold in Michigan. How will the imposition of the sugar tax impact the equilibrium price and quantity of your product?

  1. Are you happy about this new tax? Explain.
  1. If the Michigan soda lobby convinces congress not to impose this tax on Michigan-made products (i.e., still taxing Coca-Cola and Pepsi products), how do you expect this to affect the equilibrium price and quantity of your product (no explicit number needed)? Explain.
Answer #1

The demand bor faygo at the lo cal matily Jas stakon s Qp= 60- 800P and the Supply burction Cincleding Sales fae) is 6o0p-100ALSO, suppey equations Also, as=60P-1o0 Qs= -160 as=0 =) when PEo =) when 0= 600P-10 p_f0.166 e Can now driaw the demand andAnd demand eg uaton will be.- Q=600-800p =) 600.Q= gop > O.75- 0. 00125 Q =P - 72) Now. equating equation (4) and (2V we qetHnce now the equi li briem quantity will be 169.9 units aud equilibiompuće will be and pice t hat producu s receive is $ 045